OpenStudy (dawnr):

Help people! The coefficient of x^24 in the advanced level of binomial form (x^2- 2x)^13 is?

OpenStudy (dawnr):

@yamyam70

OpenStudy (yamyam70):

wait up bro ..

OpenStudy (dawnr):

waiting! :D

OpenStudy (yamyam70):



OpenStudy (dawnr):

okay

OpenStudy (yamyam70):

x^24 man is the given problem correct? I can't work this out. x^24 alone has a coefficient 1. I don't get the part of the binomial

OpenStudy (dawnr):

it should be expanded somehow

OpenStudy (dawnr):

idk xD the correct answer should be 312 but i don't know how to get there

OpenStudy (yamyam70):

okay lets expand ,

OpenStudy (yamyam70):



OpenStudy (dawnr):

lovely :D

OpenStudy (yamyam70):

well a great source of anxiety, its the perfect example. Goodluck on this one :D

OpenStudy (dawnr):

ahaha (Y) :D anxiety <3

OpenStudy (dawnr):

@ShadowLegendX

OpenStudy (dawnr):

@razor99 @tanya123 @SarahEZZMcK @Jim766 @Somy can someone help?! :O

OpenStudy (somy):

sorry T_T idk

OpenStudy (dawnr):

:/ i can't anyone who can solve this x_0

OpenStudy (dawnr):

if you know someone tag them here..

OpenStudy (dawnr):

@ganeshie8

ganeshie8 (ganeshie8):

can you write the nth term of binomial expansion ?

OpenStudy (dawnr):

..no :|

ganeshie8 (ganeshie8):

knw the binomial theorem ?

ganeshie8 (ganeshie8):

\[\large (a+b)^n = \sum \limits_{k=1}^n \binom{n}{k} a^{n-k}b^k\]

ganeshie8 (ganeshie8):

seen that before, right ?

OpenStudy (dawnr):

yea i did

ganeshie8 (ganeshie8):

good, express the given binomial power as above ^

ganeshie8 (ganeshie8):

\(\large (x^2- 2x)^{13}\) \(a = x^2\) \(b = -2x\) \(n = 13\)

ganeshie8 (ganeshie8):

plugin

ganeshie8 (ganeshie8):

\[\large (x^2 -2x)^{13} = \sum \limits_{k=1}^{13} \binom{13}{k} (x^2)^{13-k}(-2x)^k\]

ganeshie8 (ganeshie8):

\[\large ~~~ = \sum \limits_{k=1}^{13} \binom{13}{k} (x)^{26-2k}(x)^k(-2)^k\]

ganeshie8 (ganeshie8):

\[\large ~~~ = \sum \limits_{k=1}^{13} \binom{13}{k} (x)^{26-2k+k}(-2)^k\]

ganeshie8 (ganeshie8):

\[\large ~~~ = \sum \limits_{k=1}^{13} \binom{13}{k} (x)^{26-k}(-2)^k\]

ganeshie8 (ganeshie8):

Alright, looks like we're ready, what are we trying to find here ?

OpenStudy (dawnr):

the coefficient of x^24 sorry i didn't apply right away i had to go to the market

ganeshie8 (ganeshie8):

thats okay, so when do we get the x^24 ?

ganeshie8 (ganeshie8):

26 - k = 24 solve k

OpenStudy (dawnr):

the correct answer should be 312 o.O

ganeshie8 (ganeshie8):

we're not done yet, solve k first

OpenStudy (dawnr):

k=2

ganeshie8 (ganeshie8):

plugin k=2 in the formula and find the coefficient

ganeshie8 (ganeshie8):

k=2 that means the second term has a degree of 24 : \[\large ~~~ = \binom{13}{2} (x)^{26-2}(-2)^2 \]

ganeshie8 (ganeshie8):

simplify

OpenStudy (dawnr):

so i got \[\left(\begin{matrix}13 \\ 2\end{matrix}\right) (x)^{24} (4) and \left(\begin{matrix}13 \\ 2\end{matrix}\right)=78\]

ganeshie8 (ganeshie8):

multiply it by 4

OpenStudy (dawnr):

312x^24 :)

ganeshie8 (ganeshie8):

Yes !